The sum of the ages of two sisters is 32. The difference of their ages is 4. Find the age of each.
thanks so much for the help!

Answer by CharStar(110)

x = 1st age
x-4 = 2nd age
x+x-4 = 32
Add 4 to both sides add the x
2x=36
x=18
Check you answer
x=18
x-4 = 14
18+14 = 32
Hope that helps you